### v1

# Dielectronic recombination studies of ions relevant to kilonovae and non-LTE plasma

Abstract
This study presents calculations of rate coefficients, resonance strengths, and cross sections for the dielectronic recombination (DR) of $\text{Y}^{+}$, $\text{Sr}^{+}$, $\text{Te}^{2+}$, and $\text{Ce}^{2+}$—low-charge ions relevant to kilonovae and non-local thermodynamic equilibrium (non-LTE) plasmas. Using relativistic atomic structure methods, we computed DR rate coefficients under conditions typical of these environments. Our results highlight the critical role of low-lying DR resonances in shaping rate coefficients at kilonova temperatures ($\sim 10^{4}$ K) and regulating charge-state distributions. Pronounced near-threshold DR resonances significantly influence the evolving ionization states and opacity of neutron star merger ejecta. Comparisons with previous studies emphasize the necessity of including high-$n$ Rydberg states for accurate DR rate coefficients, especially for complex heavy ions with dense energy levels. Discrepancies with existing datasets underscore the need for refined computational techniques to minimize uncertainties. These results provide essential input for interpreting spectroscopic observations of neutron star mergers, including James Webb Space Telescope data. We also put forward suitable candidates for experimental studies, recognizing the challenges involved in such measurements. The data presented here have potential to refine models of heavy-element nucleosynthesis, enhance plasma simulation accuracy, and improve non-LTE plasma modeling in astrophysical and laboratory settings.
I Introduction
Dielectronic recombination (DR), i.e. the capture of a free electron via a simultaneous excitation of a bound electron followed by radiative stabilization is an atomic process that affects the ionization balance, radiative cooling, and energy distribution in both astrophysical and laboratory plasmas, particularly under non-local thermodynamic equilibrium (non-LTE) conditions [1, 2, 3, 4, 5, 6, 7, 8, 9]. In non-LTE conditions, where ionization states are not governed by local thermodynamic equilibrium, the interplay between ionization and recombination processes like DR is critical for determining plasma dynamics. This importance extends across a wide range of astrophysical and laboratory plasmas, particularly in systems characterized by diverse ionization states. An accurate treatment of DR is crucial for improving plasma models and interpreting observational data.
In laboratory plasmas, DR significantly influences energy loss mechanisms, contributing to radiative cooling and the stability of high-temperature plasmas, such as those in fusion research. The sensitivity of high Rydberg states to weak electric fields enhances DR rates even at low plasma densities, making it relevant to environments like the solar corona and magnetic confinement fusion devices [4]. By facilitating energy release through photon emission, DR aids in regulating plasma temperatures and energy distribution, critical factors for both experimental and theoretical studies of plasma behavior.
Furthermore, DR is integral to the development of collisional-radiative models, which describe the interactions between charged particles and photons in non-LTE plasmas [10, 11]. These models are essential for predicting the response of plasmas to varying conditions, including the effects of external fields that can modify ionization potentials and recombination rates. By incorporating DR, these models provide insights into phenomena such as Auger electron heating, hot electron instabilities, and ionization potential depression, which are vital for understanding plasma diagnostics and advancing non-equilibrium plasma theory [12, 13, 14].
While DR has been extensively studied in various contexts, including atomic structure and isotope shifts [15, 16, 17], its impact on kilonova-relevant ions remains an area of active investigation. Astrophysical plasmas, particularly those found in kilonovae, present extreme environments where non-LTE conditions govern the ionization structure and radiative processes. Kilonovae are bright, short-lived transients that occur following the merger of neutron stars, producing heavy elements through rapid neutron-capture (r-process) nucleosynthesis. These explosions eject high-energy material into space, evolving under extreme temperatures and densities, making them ideal laboratories for studying non-LTE plasmas. Within such environments, DR plays a dominant role in electron-ion recombination, significantly influencing ionization balance and spectral evolution. Due to its involvement with high Rydberg states, DR is highly sensitive to density effects, which govern redistributive collisions before radiative stabilization occurs. In kilonovae, which serve as key sites for rapid neutron-capture (r-process) nucleosynthesis at temperatures of approximately $10^{4}$ K, DR often competes with or even surpasses radiative recombination (RR) in shaping the ionization structure of the ejecta. Under such extreme conditions, DR becomes the predominant recombination mechanism for many heavy ions, surpassing RR [18]. This highlights the critical role of DR in shaping the ionization structure of kilonova ejecta, influencing both transient emission and the synthesis of heavy elements, mostly elements that are much heavier than Fe [19, 20, 21].
As the kilonova evolves from the diffusion phase to the nebular phase, recombination alters ionization fractions, leading to changes in spectral line intensities and the emergence or suppression of features in the optical and near-infrared regions. The rate coefficients for DR in r-process elements exhibit strong temperature dependence, particularly for singly to doubly ionized species at $T\sim 10^{4}$ K. Even slight variations in ejecta conditions can lead to substantial differences in ionization balance, directly affecting kilonova spectra. In contrast, RR typically dominates in lower-temperature environments, such as those encountered during the nebular phase of supernovae. Therefore, accurate temperature-dependent DR rates are essential for reliable spectral modeling and the interpretation of transient astrophysical events.
The importance of DR in kilonovae also stems from the complex electron shell structures of heavy elements, which result in dense energy levels and numerous autoionizing states. These states enhance the efficiency of DR. The photon emission in DR during radiative stabilization, known as dielectronic satellite emission, serves as a crucial diagnostic tool for interpreting kilonova spectra and constraining plasma parameters such as temperature and electron density. As a result, DR-driven emission plays a vital role in X-ray spectroscopy and plasma diagnostics, offering valuable insights into the physical conditions of astrophysical plasmas [4, 22].
Low-charged heavy ions such as $\text{Y}^{+},\text{Sr}^{+},\text{Te}^{2+},\text{Ce}^{2+}$ [23, 24, 25] have strong relevance to kilonovae. Observations by the James Webb Space Telescope have detected doubly charged tellurium in GRB 230307A, while singly charged strontium was detected in kilonova AT2017gfo, highlighting the importance of accurate recombination data for understanding low-charge ions in astrophysical plasmas [24, 23]. Heavy ions with excited states near the ground state (e.g., Y+, Sr+, and Ce2+) exhibit complex DR behavior even at low densities, leading to discrepancies in ionization balance calculations. Additionally, different elements contribute uniquely to kilonova opacity and spectral features; Sr and Y produce strong optical and near-infrared transitions, making them detectable in kilonova spectra, while Te and Ce, with their intricate electronic structures, significantly affect mid-infrared opacities. The recombination rates of these elements govern their ionization fractions over time, directly shaping the spectral evolution of kilonovae. Thus, incorporating precise recombination data is essential for improving kilonova models and accurately interpreting their spectral signatures.
Accurate data on DR rates, oscillator strengths, and cross-sections for electron impact are crucial for quantitative analyses of astrophysical spectra, particularly in kilonova ejecta and non-LTE plasmas. However, the scarcity of DR data for key r-process elements such as Y, Sr, Te, and Ce introduces uncertainties in modeling ionization balance and spectral evolution. The dense energy level structures of heavy elements further complicate direct comparisons of recombination rates, as these rates can vary significantly across different ions and even among transitions within the same species. Due to the experimental challenges in measuring DR rates for low-charged heavy elements, most astrophysical models rely on extrapolations from lighter elements [26] or simplified approximations, which can introduce significant inaccuracies in simulations. In particular, models of kilonovae and non-LTE plasmas often depend on crude estimates for low-charged heavy ions, where discrepancies from actual values can reach an order of magnitude, directly impacting astrophysical predictions. It is also important to note that the theoretical study of low-charge heavy ions is rather challenging. While numerous studies have focused on highly charged heavy ions [27, 28, 29, 30, 31, 32, 33, 34, 35], research on low-charge heavy ions remains very limited. To address this gap, the present work employs state-of-the-art atomic structure calculations to provide reliable DR rate coefficients for such ions. These theoretical benchmarks are essential for improving the accuracy of simulation models, refining ionization balance calculations, and enhancing the interpretation of observational data from neutron star mergers and other transient astrophysical events.
To the best of our knowledge, this work represents one of the first computations for the present targets. No experimental measurements have been found in the literature, and the only theoretical calculations available are from a recent study by Banerjee et al. [36], which covers only a few of the ions studied in this work. Another key objective is to provide reference data for experimentalists to facilitate state-of-the-art research in this field. Recent experimental advancements have enabled the study of low-charge heavy ions at the cryogenic storage ring of the Max Planck Institute for Nuclear Physics in Heidelberg [37]. Notably, successful experiments have been conducted on Xe3 ions, marking one of the highest mass-to-charge ratio DR experiments ever performed. However, such experimental efforts face numerous challenges, including the difficulty of generating such ion sources in the laboratory, managing high background noise, and detecting weak DR signals, which are often lost in the background [37]. Given the current limitations of laboratory experiments, theoretical calculations such as those presented here are required. They provide critical insights and benchmarks not only for experimentalists striving to overcome these challenges but also for astrophysicists aiming to model and interpret complex astrophysical phenomena with greater accuracy.
Section II of this article covers the theoretical specifics that are employed in the present work. The computational results are discussed and presented in Section III, while Section IV contains some concluding remarks.
II Theoretical Calculations
The three DR reactions studied in the present work are:
(i) for Y II to Y I, (for $l\leq 9$ and $n=6,\dots,35$)
${\rm Y}^{+}[5s^{2}(i)]+e^{-}$$\to\left\{\begin{array}[]{l}{\rm Y}^{*}[5s~{}5p~{}nl_{j}(d)]\\ {\rm Y}^{*}[5s~{}4d~{}nl_{j}(d)]\end{array}\right\}\to{\rm Y}(f)+h\nu\,,$ (3)
(ii) for Sr II to Sr I, (for $l\leq 9$ and $n=6,\dots,35$)
${\rm Sr}^{+}[5s(i)]+e^{-}$$\to\left\{\begin{array}[]{l}{\rm Sr}^{*}[5p~{}nl_{j}(d)]\\ {\rm Sr}^{*}[4d~{}nl_{j}(d)]\end{array}\right\}\to{\rm Sr}(f)+h\nu\,,$ (6)
(iii) for Te III to Te II, (for $l\leq 9$ and $n=6,\dots,35$)
${\rm Te}^{2+}[5s^{2}~{}5p^{2}(i)]+e^{-}$$\to\left\{\begin{array}[]{l}{\rm Te}^{+*}[5s^{2}~{}5p~{}5d~{}nl_{j}(d)]\\ {\rm Te}^{+*}[5s~{}5p^{3}~{}nl_{j}(d)]\end{array}\right\}$ (9)
$\to{\rm Te}^{+}(f)+h\nu\,,$ (10)
(iv) for Ce III to Ce II, (for $l\leq 9$ and $n=6,\dots,24$)
${\rm Ce}^{2+}[4f^{2}(i)]+e^{-}$$\to\left\{\begin{array}[]{l}{\rm Ce}^{+^{*}}[4f~{}5d~{}nl_{j}(d)]\end{array}\right\}$ (12)
$\to{\rm Ce}^{+}(f)+h\nu\,.$ (13)
Here, $i$ denotes the ground state, $d$ represent the intermediate states, and $f$ indexes the final states. The Rydberg electron is represented as $nl_{j}$, where $n$ is the principal quantum number of the captured electron. Here, $n$ is determined based on convergence. Including more $n$-values (e.g., in the hundreds) would have some minor effect on the results, but the difference should not be very significant compared to the values used in the present case. This contributes to some uncertainty in the calculation, which has been accounted for when addressing the overall uncertainty. The emitted decay photons are denoted by $h\nu$. For a given $n$, angular momentum states $l=0,1,\dots,9$ and $j=|l\pm\frac{1}{2}|$ are included. Radiative decay involves all electric dipole transitions to lower-lying states resulting in a significant number of states to consider. Except for the $4d$ state of Sr+ and $4f$ state of Ce2+, all transitions occur within the same shell ($\Delta n=0$), while the $5s\to 4d$ transition in Sr+ and $4f\to 5d$ transition in Ce2+ involves inter-shell excitation ($\Delta n=1$).
For a dielectronic recombination channel, i.e. for a two-step transition $i\to d\to f$, the cross section is expressed as a function of the electron kinetic energy $E$ in the independent resonances approximation as (see, e.g. [38, 39, 40, 41, 32])
$\sigma^{{\mathrm{DR}}}_{i\to d\to f}(E)=\frac{2\pi^{2}}{p^{2}}V_{a}^{i\to d}\frac{A_{r}^{d\to f}}{\Gamma_{d}}L_{d}(E).$ (14)
The initial state of the DR process which consists of the ground-state ion and a continuum electron with an asymptotic momentum $\vec{p}$ and spin projection $m_{s}$. In addition, $\Gamma_{d}$ is the total natural width of the intermediate autoionizing state, which is the sum of the radiative and autoionization widths: $\Gamma_{d}=A^{d}_{r}+A^{d}_{a}$ (here in atomic units with $\hbar=1$). $L_{d}(E)$ is the Lorentzian line shape function, expressed as
$L_{d}(E)=\frac{\Gamma_{d}/(2\pi)}{(E_{i}+E-E_{d})^{2}+\frac{\Gamma_{d}^{2}}{4}},$ (15)
and is normalized to unity on the energy scale where $p=|\vec{p}|=\sqrt{(E/c)^{2}-c^{2}}$ is the modulus of the free-electron momentum associated with the kinetic energy $E$.
The dielectronic capture rate is related to the rate of its time-reversed process, i.e., the Auger process, by the principle of detailed balance:
$V_{a}^{i\to d}=\frac{2J_{d}+1}{2(2J_{i}+1)}A_{a}^{i\to d}\,.$ (16)
Here, $J_{d}$ and $J_{i}$ are the total angular momenta of the intermediate and the initial states of the recombination process, respectively. Neglecting the energy-dependence of the electron momentum in the vicinity of the resonance, the dielectronic resonance strength, defined as the integrated cross section for a given resonance peak,
[…]
is given as
$S^{\mathrm{DR}}_{i\to d\to f}=\frac{2\pi^{2}}{p^{2}}\frac{1}{2}\frac{2J_{d}+1}{2J_{i}+1}\frac{A_{a}^{i\to d}A_{r}^{d\to f}}{A_{r}^{d}+A_{a}^{d}}\,,$ (18)
where $A_{a}^{i\to d}$ is implicitly defined in Eq. (16). The factor $\frac{2\pi^{2}}{p^{2}}$ defines the phase space density and the $1/2$ stems from the spin degeneracy of the free electron.
The total rate coefficients ($\alpha_{DR}$) for astrophysical and thermal plasmas are described by
$\alpha_{DR}(T)$ $=\frac{h^{3}}{(2\pi m_{e}kT)^{3/2}}\sum_{d}\frac{2J_{d}+1}{2(2J_{i}+1)}$
×$\quad\times\frac{A_{a}^{i\to d}A_{r}^{d\to f}}{A_{r}^{d}+A_{a}^{d}}\exp\left(-\frac{E}{kT}\right)\,,$ (19)
derived by summing across all possible autoionization channels and averaging over the Maxwellian distribution of electron energies [42]. In this expression, $k$ denotes the Boltzmann constant, $h$ is the Planck constant that we write explicitly here, $T$ represents the electron temperature, and $E$ is the resonance energy.
III Results and Discussion
In this study, the relativistic configuration interaction method with independent-particle basis wave functions was utilized to calculate the energy levels, radiative rates, autoionization rates, and DR cross sections and rates as implemented in the Flexible Atomic Code (FAC) [43, 44, 45]. The relativistic distorted-wave approximation was employed to describe the continuum states. For the calculation of wave functions and energy levels corresponding to the initial states, intermediate doubly excited states, and radiative final states, contributions from electron correlations, quantum electrodynamics (QED) effects, and Breit interactions were systematically accounted for, ensuring accurate determination of energy levels and wave functions.
Figure 1: Total DR cross section for Y+ recombining into Y is plotted against the relative electron energy. The Lorentzian line shapes are convoluted with a Gaussian function with a width of 100-meV.
Figure 1 illustrates the total DR cross sections for the recombination of Y+ into Y. The resolution width is set to 100 meV, modeled as a Lorentzian profile convoluted with a Gaussian. The calculations include Rydberg states up to $n=6,\ldots,35$. A strong DR resonance is observed near the threshold, accompanied by relatively weaker resonance peaks at approximately 0.3 eV and 0.6 eV. These peaks are attributed to the $4d5s~{}^{3}D_{J}$ states (where $J=1,2,3$). Additionally, a small resonance around 2 eV is associated with the $5s5p~{}^{3}P_{J}$ states (where $J=0,1,2$).
The DR resonance strength for Y+ is approximately an order of magnitude higher than that of previous ions studied by our group [45, 37]. The presence of strong resonances at low energies suggests that DR is dominated by optically allowed core excitations, and it is a dominant recombination mechanism for Y+ under various plasma conditions, particularly in astrophysical environments. Given the exceptionally strong resonance near the threshold, Y+ is an excellent candidate for experimental validation, as background effects are expected to be minimal, enhancing detection feasibility. However, experimental challenges remain, particularly in distinguishing the DR signal from radiative recombination, which is typically prominent at near-threshold energies. Effectively separating these contributions will be crucial for future measurements. The findings of this work could facilitate both observational and experimental studies of Y+ in the near future.
Figure 2: Total DR cross section for Sr+ recombining into Sr is plotted against the relative electron energy.
The total DR spectrum for Sr+ recombining to neutral Sr is presented in Figure 2, with a resolution width set to 100 meV. The DR cross section exhibits characteristic resonance peaks, arising from the capture of free electrons into autoionizing Rydberg states via core excitations ($5s\to 5p,4d$). A strong DR resonance feature is observed near zero energy; however, its exact origin remains unidentified. There is a strong possibility that it results from a theoretical artifact associated with the Rydberg state $n=6$, as observed in our calculations, since no such excited states have been reported in the literature at such low energies. Another prominent resonance appears at approximately 2.3 eV, primarily originating from the $4p^{6}4d~{}^{2}D_{5/2}$ and ${}^{2}D_{3/2}$ states, with the ${}^{2}D_{5/2}$ state being the dominant contributor. Additionally, a strong DR resonance is observed around 2.7 eV, attributed to the $4p^{6}5p~{}^{2}P_{1/2}$ and ${}^{2}P_{3/2}$ states. These states are also responsible for the knee-like resonances observed around 3 eV.
Among the ions studied in this work, the DR spectrum of Sr+ exhibits the weakest signal strength, possibly presenting significant challenges for experimental observation with currently available setups. Its DR signal strength is approximately one to two orders of magnitude smaller than that of the other ions studied in this work as well as in previous works[45, 37]. To detect this spectrum experimentally, it would be necessary to significantly reduce background noise.
Figure 3: Total DR cross section for Te2+ recombining into Te+ is plotted against the relative electron energy.
Figure 3 presents the total DR cross section for Te2+ recombining into Te+, calculated with a resolution width of 100 meV. In the DR spectrum, a small resonance is observed at low energies, below 0.5 eV, attributed to the $5s^{2}5p^{2}~{}^{3}P_{J}$ states (where $J=0,1,2$). At approximately 14.5 eV, two prominent resonance structures appear, corresponding to the $5p5d~{}^{3}D_{J}$ (where $J=1,2,3$) and ${}^{3}P_{J}$ (where $J=0,1,2$) states, with the ${}^{3}D_{1}$ and ${}^{3}P_{1}$ states being the dominant contributors. Additionally, several significant resonance peaks are observed on either side of this strong resonance feature, arising from the $5s5p^{3}$ states.
Given the presence of a well-defined resonance near the threshold and strong DR resonance features at higher energies, along with a relatively high DR signal strength, this ion is a promising candidate for further DR spectroscopy studies, both experimentally and theoretically.
Figure 4: Total DR cross section for Ce2+ recombining into Ce+ is plotted against the relative electron energy.
Figure 4 presents the total dielectronic recombination (DR) cross sections for the recombination of Ce2+ into Ce+, calculated with a resolution width of 100 meV. A strong DR resonance is observed near the threshold, primarily due to the convolution of the $4f^{2}$ and $4f~{}5d$ states. At energies close to 0 eV, the $4f^{2}~{}^{3}H_{5}$ state is the dominant contributor, whereas toward the tail end of the peak, the $4f~{}5d~{}^{1}G_{4}$ state becomes the most significant. Although the spectrum appears to decline beyond 0.2 eV, the cross-section magnitude remains on the order of $10^{7}$ barns, emphasizing the strength of this prominent resonance feature. This suggests that Ce2+ could serve as a strong candidate for experimental study near the threshold energy, similar to Y+. Furthermore, the exceptionally high DR signal might indicate a considerable probability of its detection in a kilonova.
The findings of Gu [43] suggest that the energy values computed using the FAC may have uncertainties due to factors such as incomplete treatment of correlations or limitations in the atomic model. Furthermore, according to the FAC manual authored by Gu [46], in the case of near-neutral ions (as in the present case), the uncertainties associated with radiative decay rates and autoionization rates can reach 20% or, in certain instances, exceed this value. Additionally, limiting the number of states $n$ introduces an uncertainty of approximately 10%. Therefore, the overall minimum uncertainty in the present calculation of DR cross section and strengths is estimated to be around 30%. Given the substantial amount of data generated in this work, all data are made available as supplementary material.
Figure 5: Rate-coefficient for Y+, Sr+, Te2+, and Ce2+ is plotted against the electron temperature.
Figure 5 illustrates the dependence of the total DR rate coefficients on electron temperature, calculated by incorporating Rydberg states with principal quantum numbers ranging from $n=6$ to $n=35$, except for $\text{Ce}^{2+}$. While the high-temperature rates ($T>10^{5}$ K) may not have significant practical relevance, they are included for completeness. At high temperatures $T$, the DR rate coefficient decreases following a $\sim T^{-3/2}$ dependence, as the increasing kinetic energy of free electrons reduces the probability of recombination. In contrast, at low temperatures, the DR process is predominantly influenced by a number of resonances situated just above the threshold. These resonances contribute significantly to the recombination rate, ensuring that DR remains efficient even at lower temperatures. This occurs because the $\sim T^{-3/2}$ factor compensates for the lowness of the temperature. The calculated values have an estimated uncertainty of at least 38%, determined using the error propagation rule from Eq. II. At low temperatures, DR is primarily governed by near-zero-energy resonances, where the energy of free electrons closely matches the energy levels of states within the ion, enabling highly efficient recombination processes. The dominance of these resonances enhances the recombination probability, significantly influencing the ionization balance of the kilonova ejecta. Notably, the behavior of DR at these temperatures can resemble radiative recombination, highlighting the need for precise resonance data for accurate modeling. As the temperature increases, contributions from high-$n$ Rydberg states become more prominent. These states, with their highly excited electrons near the ionization threshold, play a crucial role in recombination dynamics at higher temperatures. The interplay between the thermal energy of the electrons and the populations of these excited states alters the overall recombination efficiency, resulting in a shift in the effective ionization balance. This shift is particularly significant for ions with one or two electrons outside closed shells (e.g., $\text{Y}^{+}$, $\text{Sr}^{+}$, and $\text{Ce}^{2+}$), as it impacts the interpretation of kilonova spectra and the modeling of non-LTE plasmas [7].
Table 1: DR rate-coefficients at temperature $10^{4}$ K.
Ion Rate-coefficient ($10^{-11}\,\text{cm}^{3}/\text{s}$)
Y+ 9.61403
Sr+ 0.21363
Te2+ 0.31646
Ce2+ 10.96778
Among the studied ions, $\text{Ce}^{2+}$ exhibits relatively high DR rate coefficients at very low temperatures compared to the other ions. In the kilonova-relevant temperature range ($\sim 10^{4}$ K), as shown in Table 1, the rate coefficients for $\text{Y}^{+}$ and $\text{Ce}^{2+}$ become comparable and significantly exceed those of $\text{Sr}^{+}$ and $\text{Te}^{2+}$. This indicates that $\text{Y}^{+}$ and $\text{Ce}^{2+}$ have a stronger tendency to populate lower ionization states over time, while $\text{Sr}^{+}$ and $\text{Te}^{2+}$ remain ionized for a longer duration in the expanding ejecta. These DR trends have critical implications for kilonova spectra. Higher DR rates, as seen for $\text{Y}^{+}$ and $\text{Ce}^{2+}$, are expected to promote efficient recombination to lower ionization states, contributing to optical and near-infrared spectral features. In contrast, the slower recombination associated with $\text{Sr}^{+}$ and $\text{Te}^{2+}$ can lead to the retention of higher ionization states, producing spectral features characterized by higher-energy transitions and shifts toward shorter wavelengths. The availability of these computed DR rates will be instrumental in refining spectral models, improving predictions of line emission, and enhancing our understanding of kilonova ejecta composition, as they provide crucial inputs for astrophysical modeling tools like the SUMO spectral synthesis code used in kilonovae simulations [47].
Figure 6: Comparative study for rate-coefficient for Y+, and Sr+ with available study from Banerjee et al. [36].
In Figure 6, a comparison is made between the DR rate coefficients obtained in this work and those reported in the only existing study over the temperature range characteristic of the nebular phase of kilonovae. During this phase, the temperature is typically around 10,000 K [48, 49]. To account for the temperature evolution from the early weeks after the event to later stages, influenced by the evolving ejecta composition, the rate coefficients are presented across a broader temperature range from 1,000 K to 30,000 K.
In the case of $\text{Y}^{+}$, the present values are approximately an order of magnitude larger than those reported by Banerjee et al. [36]. The same number of intermediate states has been considered in both studies; however, the difference in magnitude arises from the inclusion of a significantly larger number of Rydberg states in the present work. The inclusion of high-$n$ Rydberg states enhances the recombination probability [50]. Additionally, differences in the number of orbital angular momentum states ($l$) considered in the two studies may further contribute to the observed discrepancy in the rate coefficients.
A different trend is observed for $\text{Sr}^{+}$, where the present results show good agreement with those of Banerjee et al. [36]. This consistency can be attributed to the inclusion of an additional $4p^{6}4f$ state in their study, which increases their rate coefficient. However, this enhancement is counterbalanced by their lower number of Rydberg states and $l$-values, resulting in an overall accidental numerical agreement between the two datasets. The role of Rydberg states is particularly significant in singly charged ions like $\text{Y}^{+}$ and $\text{Sr}^{+}$, where dense energy levels allow for extensive recombination pathways [51]. Despite some variations in magnitude, the present DR rate coefficients remain within an order of magnitude—or better—compared to previous studies. This level of consistency supports the reliability of the present work.
IV Conclusions
Given the importance of DR in kilonova spectra and non-LTE plasmas, systematic calculations of DR rates, strengths, and cross sections for low-charge heavy ions such as $\text{Y}^{+}$, $\text{Sr}^{+}$, $\text{Te}^{2+}$, and $\text{Ce}^{2+}$ are critical. These data are essential for interpreting transient emissions across multiple electromagnetic bands in kilonovae and improving plasma models for astrophysical and laboratory contexts. Studying DR processes in non-LTE plasmas enhances our understanding of plasma behavior across diverse environments, advancing theoretical and experimental plasma physics.
Strong near-threshold DR resonances for $\text{Y}^{+}$ and $\text{Ce}^{2+}$ highlight their role in modifying the ionization balance of kilonova ejecta. DR rate sensitivity to plasma conditions, such as temperature and density, underscores the need for precise data to refine ionization state modeling. These rate coefficients are crucial for interpreting observational data from instruments like the James Webb Space Telescope, as seen in the detection of tellurium in specific kilonovae. Additionally, the methodology employed in this study can be extended to further DR investigations, improving simulations beyond crude approximations currently in practice.
Comparing our calculated DR rates with previous studies reveals the impact of including high-$n$ Rydberg states, illustrating challenges in achieving accurate results for complex heavy ions. Discrepancies between our data and other studies highlight the need for continued advancements in computational techniques to reduce uncertainties in DR rate coefficients. This work identifies several candidates for experimental measurements, such as $\text{Y}^{+}$ and $\text{Ce}^{2+}$ for near-threshold DR studies due to their strong DR signal strengths. $\text{Te}^{2+}$ also emerges as a promising candidate for DR spectrum studies and benchmarking DR analyses. These findings enhance interpretations of transient phenomena and provide a foundation for experimental validation.